3. Optimal Consumption Decision in Utility Maximization Model
1) Biwei’s marginal use value for chamomile tea consumption is MU=-2Q+6, where Q is the box
quantity of chamomile tea consumed and takes only integer values. Draw the marginal use value
curve and total use value curve for Q less than 5. If the price for a box of chamomile tea is $4, how
many boxes will Biwei consume per week? If the price drops to $2, how many boxes will Biwei
consume per week?
2) Suppose Biwei’s total use value is now given by TU = 8√Q, where Q takes only integer values.
Draw the marginal use value curve and total use value curve for Q less than 5. If the price for a
box of chamomile tea is $4, how many boxes will Biwei consume per week? If the price drops to
$2, how many boxes will Biwei consume?3. Optimal Consumption Decision in Utility Maximization Model 1) Biweis marginal use value for chamomile tea consumption is
Answer #1

1)bune HMUCuxua 12 3 4 5 гого,05,0;

The optimal consumption will be at,

P( additional unit cost)=MU( additional unit benefit)

Given p=4

So 4=6-2Q

Q=1,

Given price equal to 4 ,biwei consume 1 unit

P=2

So,2=6-2Q

Q=2

Given price Equal to 2, biwei consume 2 units.

2)TU=8√Q

MU=4/√Q

TO To my 7 1 2 3 4 5 6

​​​

The optimal consumption will be at,

P( additional unit cost)=MU( additional unit benefit)

Given p=4

So 4=4√Q

Q=1,

Given price equal to 4 ,biwei consume 1 unit

P=2

So,2=4√Q

√Q=0.5

Q=0.25

Given price Equal to 2, biwei consume 0.25 units.